# MAPF-HD: Multi-Agent Path Finding in High-Density Environments

### v1

@@ IV-D Why Does PHANS Terminate Within a Finite Number of Steps?
The proposed method prioritizes path planning for obstructing agents on the target paths by assigning a higher priority to those farthest from the goal. This definition ensures that at least one empty vertex exists in the environment. This vertex is assigned to the obstructing agent with the longest path to the goal. $v^{\ast}$ denotes the vertex of the obstructing agent.
@@
A finite-length path exists between $v^{\ast}$ and the nearest empty vertex (null agent), and null-agent swapping is performed at each time step along this path. Consequently, $d^{t}_{\mathrm{null}}(v^{\ast})$ decreases monotonically at each time step $d_{\mathrm{null}}^{t}(v^{\ast})>d_{\mathrm{null}}^{t+1}(v^{\ast})$. This process continues until $d^{t}_{\mathrm{null}}(v^{\ast})$ reaches zero after a finite number of steps. At this point, $v^{\ast}$ becomes empty, allowing the farthest target agent to advance to a newly vacated vertex.
@@ V Numerical Experiments
This section describes the four experiments conducted to evaluate the performance of the proposed method. Experiments 1 and 2 compared PHANS with PIBT, EECBS, and ILP in a small-scale simulation environment where these methods could be applied. Experiment 3 assessed the scalability of PHANS in large environments, demonstrating its effectiveness. Additionally, Experiment 4 assessed the applicability of PHANS in complex environments with obstacles such as pillars. All experiments were conducted on an Intel Core i9-12900K processor with 128 GiB of RAM and Ubuntu 22.04.
V-A Comparison Methods Setting
@@
Unlike PHANS and ILP, PIBT and EECBS require predefined goals for agents whose paths are planned. In the experiments, we assign the obstructing agents goals that return them to their starting positions. However, the obstructing agents stop moving once all target agents reach their goals. In other words, the makespan (defined in Section III-D) is computed only for the target agents.
@@
PIBT can assign priorities to agents. Accordingly, we set priority values to the distance between target agents’ start and goal positions and assigned a priority of 0 to the obstructing agents. Because vanilla PIBT does not consider following conflicts, we extended PIBT using the method described in [23].
@@ V-B Experiment 1
The simulation environment is shown in Fig. 5. The simulations involved two target agents, where the number of obstructing agents varied from $0\%$ to $90\%$ of the total number of cells. The objective was to minimize the makespan and move the target agents to their destinations without collisions.
@@ V-B1 Success Rate and Computation Time
Compared with ILP and EECBS, PIBT required a lower computation time; specifically $0.11$ s up to at a density of $80\%$. The computation time increased considerably at a density of $90\%$, attributed to the computational complexity of PIBT. PIBT does not know where empty cells exist prior to the evacuation of obstructing agents. Agents in PIBT only know their goals, and the evacuating position (selecting empty cells) is not optimized. To avoid following conflicts, PIBT performs a random search to determine the evacuating position. If the probability that a cell is occupied is $p$, the probability of finding an empty cell in a given number of search attempts is $p^{times-1}\cdot(1-p)$, where $p$ is assumed to be fixed and independent of previous visit paths. At a density of $90\%$, the obstructing agents struggle to find the $10\%$ unoccupied cells. The computational complexity at $90\%$ density is much higher than at lower densities, which consequently increases the computation time.
@@ V-B2 Makespan
PIBT has a much larger makespan than that of PHANS. PIBT can resolve local conflicts (around the target agents) via backtracking. However, backtracking cannot be applied to future conflicts, and moving the obstructing agents in advance is challenging. Therefore, the number of local-conflict resolutions increases, which increases the makespan.
@@ V-C Experiment 2
Experiment 1 was conducted in an environment without obstacles. Experiment 2 was conducted in an environment with pillar-like obstacles (Fig. 7). We set the number of target agents to $12$, and the cell occupation density varied from $20\%$ to $90\%$ of the cells. The densities included obstructing agents (varied) and static obstacles (fixed at $90$ cells). As in Experiment 1, the number of target agents was set to $2$, and the maximum computation time for each trial was set to $180$ s.
@@ V-C1 Success Rate and Computation Time
The success rates and computation times (Figs. 8(a) and (b)) of PHANS, EECBS, and ILP were similar to those in Experiment 1. PIBT achieved shorter computation times than in Experiment 1, particularly at $90\%$ density. PIBT moved obstructing agents to empty cells via backtracking. As static obstacle cells were excluded from the search, the search space was reduced compared with environments without obstacles, ultimately resulting in a smaller computational time and improved success rate.
@@ V-C2 Makespan
The makespans (Fig. 8(c)) of EECBS and ILP were similar to those in Experiment 1, whereas PHANS and PIBT achieved shorter makespans under certain conditions. This is presumably because the presence of static obstacles reduces the search space, making the evacuation of obstructing agents to favorable locations easier.
@@ V-D1 Success Rate and Computation Time
Figs. 10(a) and (b) show the success rate and computation time, respectively. PHANS completed all path-planning tasks with minimal computation time across various densities. Similar to Experiment 1, even at the highest density ($95\%$ in Experiment 3), the average computation time remained low at $2.07$ s.
Fig. 10(b) indicates that higher obstructing agent densities generally led to longer computation time. Increasing the number of obstructing agents on the target paths requires more evacuations. This increases the number of calculations required for null-agent paths, which increases the computation time.
Fig. 10(b) shows that the computation time of PHANS decreased when the density of obstructing agents exceeded $80\%$. Even at a density of $95\%$, the average computation time was only $2.07$ s. PHANS assigns priorities to the obstructing agents, and null agents are systematically allocated until no further assignments are possible (Algorithm 1, lines 6–16). At extremely high densities, the limited availability of null agents simplifies the assignment process and reduces the computation time. By contrast, ILP faces tighter global coupling on the time-expanded network, EECBS resolves more conflicts as density rises, and PIBT incurs heavier backtracking when feasible local moves vanish. Thus, only PHANS exploits the lack of empty cells to shrink the assignment space, which lowers the computation time at high occupation density.
The success rates and computation times of PIBT, EECBS, and ILP were similar to those in Experiment 1. The computational cost of PIBT increases because, in extremely high-density environments, the probability of having an empty neighboring cell decreases, which increases the backtracking computational cost. In ILP, the costs of preprocessing and constructing the time-expanded network increase exponentially with the scale of the environment. Consequently, even when the upper limit of the maximum computation time was relaxed from 180 s to 3600 s, ILP failed to find a feasible solution.
@@ V-D2 Makespan
As in Experiment 1, the makespan of PHANS was significantly smaller than that of PIBT. In large-scale environments, the target path becomes longer, and the number of obstructing agents that need to be moved increases. At this point, moving obstructing agents in advance has a significant effect.
@@
Figure 12: Comparison of (a) success rates, (b) computation time, and (c) makespan between PHANS and PIBT in Experiment 4. The density is the proportion of cells occupied by obstructing agents and static obstacles.
@@ V-E Experiment 4
We evaluated the applicability of PHANS in complex environments with obstacles such as pillars, which are common in large warehouse environments. In this scenario, path planning is challenging. For Experiment 4 (Fig. 11), we set static obstacles as in Experiment 2 and assessed the makespan and computation time across varying densities. We set the number of target agents to 12, whereas the occupation density varied from $15\%$ to $95\%$ of the cells. The densities included obstructing agents (varied) and static obstacles (fixed at $90$ cells). Fifty trials were performed for each density. As in Experiments 1–3, the maximum computation time for each trial was $180$ s.
@@ V-E1 Success Rate and Computation Time
As in Experiment 3, PHANS demonstrated a high computational efficiency in the obstructed environment considered in Experiment 4. Compared with PIBT, paths were successfully planned in most trials with significantly shorter computation time (Figs. 12(a) and (b)). However, at a density of $95\%$, PHANS failed to plan a path in one trial. PHANS prioritizes target agents based on their distance from the goal and statically determines target paths. Owing to high-priority target agents and static obstacles, a deadlock occurred in one trial during the evacuation of obstructing agents, causing path-planning failure.
@@ VI Conclusion
This study defined MAPF-HD and introduced a new heuristic method, PHANS, to solve MAPF-HD efficiently. PHANS operates in two phases: first, it plans the path for the target agents, and then it sequentially evacuates obstructing agents to clear the path. This approach involves swapping positions between agents and empty vertices, which we refer to as null agents.
@@
However, PHANS has limitations in complex scenarios. For instance, its performance can degrade in environments with a large number of static obstacles or a high ratio of target agents; nonetheless, such conditions are less typical in automated warehouses and valet parking applications. PHANS does not explicitly handle dynamic obstacles; however, its rapid planning speed allows for frequent re-planning when such obstacles appear.
Overall, our results highlight PHANS as a scalable and practical approach for real-world applications such as high-density warehouse logistics and valet parking systems. Future research will focus on adaptive target paths and null-agent assignment strategies. Additionally, we will extend PHANS to handle more complex scenarios and general graphs.

### v2

@@ IV-D Why Does PHANS Terminate Within a Finite Number of Steps?
The proposed method prioritizes path planning for obstructing agents on the target paths by assigning a higher priority to those farthest from the goal. By definition of the environment, at least one empty vertex exists; therefore, this vertex can be assigned to the obstructing agent with the longest path to its goal. Let $v^{\ast}$ denote the vertex of the obstructing agent.
@@
A finite-length path exists between $v^{\ast}$ and the nearest empty vertex (null agent), and null-agent swapping is performed at each time step along this path. Consequently, $d^{t}_{\mathrm{null}}(v^{\ast})$ decreases monotonically at each time step, i.e., $d_{\mathrm{null}}^{t}(v^{\ast})>d_{\mathrm{null}}^{t+1}(v^{\ast})$. This process continues until $d^{t}_{\mathrm{null}}(v^{\ast})$ reaches zero after a finite number of steps. At this point, $v^{\ast}$ becomes empty, allowing the farthest target agent to advance to a newly vacated vertex.
@@ V Numerical Experiments
This section describes the four experiments conducted to evaluate the performance of the proposed method. Experiments 1 and 2 compared PHANS with PIBT, EECBS, and ILP in small-scale simulation environments where these methods could be applied. Experiment 3 assessed the scalability of PHANS in large environments, demonstrating its effectiveness. Additionally, Experiment 4 assessed the applicability of PHANS in complex environments with obstacles such as pillars. All experiments were conducted on an Intel Core i9-12900K processor with 128 GiB of RAM and Ubuntu 22.04.
V-A Settings for Comparison Methods
@@
Unlike PHANS and ILP, PIBT and EECBS require predefined goals for agents whose paths are planned. In the experiments, we assigned the obstructing agents goals that return them to their starting positions. However, the obstructing agents stopped moving once all target agents reached their goals. In other words, the makespan was computed only for the target agents.
@@
PIBT can assign priorities to agents. Accordingly, we set each target’s priority equal to its start-goal distance and assigned a priority of 0 to obstructing agents. Because vanilla PIBT does not consider following conflicts, we extended it using the method described in [23].
@@ V-B Experiment 1
The simulation environment is shown in Fig. 5. The simulations involved two target agents, and the number of obstructing agents varied from $0\%$ to $90\%$ of the total number of cells. The objective was to minimize the makespan and move the target agents to their destinations without collisions.
@@ V-B1 Success Rate and Computation Time
Compared with ILP and EECBS, PIBT required less computation time; at a density of $80\%$, it took only $0.11$ s. The computation time increased considerably at a density of $90\%$, due to the computational complexity of PIBT. PIBT does not know where empty cells exist prior to the evacuation of obstructing agents. Agents in PIBT only know their goals, and the evacuating position (selecting empty cells) is not optimized. To avoid following conflicts, PIBT performs a random search to determine the evacuating position. If the probability that a cell is occupied is $p$, the probability of finding an empty cell in a given number of search attempts is $p^{\mathrm{times}-1}\cdot(1-p)$, where $p$ is assumed to be fixed and independent of previously visited paths. At a density of $90\%$, the obstructing agents struggle to find the $10\%$ unoccupied cells. The computational complexity at $90\%$ density is much higher than at lower densities, which consequently increases the computation time.
@@ V-B2 Makespan
PIBT has a much larger makespan than PHANS. PIBT can resolve local conflicts (around the target agents) via backtracking. However, backtracking cannot be applied to future conflicts, and moving the obstructing agents in advance is challenging. Therefore, the number of local-conflict resolutions increases, which increases the makespan.
@@ V-C Experiment 2
Experiment 1 was conducted in an environment without obstacles. Experiment 2 was conducted in an environment with pillar-like obstacles (Fig. 7). We set the number of target agents to $2$ and the cell occupation density varied from $20\%$ to $90\%$ of the cells. The densities included obstructing agents (varied) and static obstacles (fixed at $90$ cells). As in Experiment 1, the maximum computation time for each trial was set to $180$ s.
@@ V-C1 Success Rate and Computation Time
The success rates and computation times (Figs. 8(a) and (b)) of PHANS, EECBS, and ILP were similar to those in Experiment 1. PIBT achieved shorter computation times than in Experiment 1, particularly at $90\%$ density. PIBT moved obstructing agents to empty cells via backtracking. As static obstacle cells were excluded from the search, the search space was reduced compared with environments without obstacles, ultimately resulting in a smaller computation time and improved success rate.
@@ V-C2 Makespan
The makespans (Fig. 8(c)) of EECBS and ILP were similar to those in Experiment 1, whereas PHANS and PIBT achieved shorter makespans under certain conditions. This is presumably because the presence of static obstacles reduces the search space, making it easier to evacuate obstructing agents to favorable locations.
@@ V-D Experiment 3
Figure 12: Comparison of (a) success rates, (b) computation time, and (c) makespans between PHANS and PIBT in Experiment 4. The density is the proportion of cells occupied by obstructing agents and static obstacles.
@@ V-D1 Success Rate and Computation Time
Figs. 10(a) and (b) show the success rate and computation time, respectively. PHANS completed all path-planning tasks with minimal computation time across various densities. Similar to Experiment 1, even at the highest density ($95\%$), the average computation time remained low at $2.07$ s.
Fig. 10(b) indicates that higher obstructing agent densities generally led to longer computation times. Increasing the number of obstructing agents on the target paths requires more evacuations. This increases the number of calculations required for null-agent paths, which increases the computation time.
Fig. 10(b) shows that the computation time of PHANS decreased when the density of obstructing agents exceeded $80\%$. Even at a density of $95\%$, the average computation time was only $2.07$ s. PHANS assigns priorities to the obstructing agents, and null agents are systematically allocated until no further assignments are possible (Algorithm 1, lines 6–16). At extremely high densities, the limited availability of null agents simplifies the assignment process and reduces the computation time. By contrast, ILP faces tighter global coupling on the time-expanded network; EECBS resolves more conflicts as density rises; and PIBT incurs heavier backtracking when feasible local moves vanish. Thus, only PHANS exploits the lack of empty cells to shrink the assignment space, which lowers the computation time at high occupation density.
The success rates and computation times of PIBT, EECBS, and ILP were similar to those in Experiment 1. The computational cost of PIBT increases because, in extremely high-density environments, the probability of having an empty neighboring cell decreases, which increases the backtracking computational cost. In ILP, the costs of preprocessing and constructing the time-expanded network increase exponentially with the scale of the environment. Consequently, even when the computation time limit was relaxed from 180 s to 3600 s, ILP failed to find a feasible solution.
@@ V-D2 Makespan
As in Experiment 1, the makespan of PHANS was significantly smaller than that of PIBT. In large-scale environments, target paths become longer, and the number of obstructing agents that need to be moved increases. At this point, moving the obstructing agents in advance has a significant effect.
@@ V-E Experiment 4
We evaluated the applicability of PHANS in complex environments with obstacles such as pillars, which are common in large warehouse environments. For Experiment 4 (Fig. 11), we set static obstacles as in Experiment 2 and assessed the makespan and computation time across varying densities. We set the number of target agents to 12, whereas the occupation density varied from $15\%$ to $95\%$ of the cells. The densities included obstructing agents (varied) and static obstacles (fixed at $90$ cells). Fifty trials were performed for each density. As in Experiments 1–3, the maximum computation time for each trial was $180$ s.
@@ V-E1 Success Rate and Computation Time
As in Experiment 3, PHANS demonstrated high computational efficiency in the obstructed environment considered in Experiment 4. Compared with PIBT, PHANS successfully planned paths in most trials with significantly shorter computation time (Figs. 12(a) and (b)). However, at a density of $95\%$, PHANS failed to plan a path in one trial. PHANS prioritizes target agents based on their distance from the goal and statically determines target paths. Owing to high-priority target agents and static obstacles, a deadlock occurred in one trial during the evacuation of obstructing agents, causing path-planning failure.
@@ VI Conclusion
This study defines MAPF-HD and introduces a new heuristic method, PHANS, to solve MAPF-HD efficiently. PHANS operates in two phases: first, it plans paths for the target agents, and then it sequentially evacuates obstructing agents to clear the path. This approach involves swapping positions between agents and empty vertices, which we refer to as null agents.
@@
However, PHANS has limitations in certain complex scenarios. For instance, its performance can degrade in environments with a large number of static obstacles or a high ratio of target agents; nonetheless, such conditions are less typical in automated warehouses and valet parking applications. PHANS does not explicitly handle dynamic obstacles; however, its rapid planning speed allows for frequent re-planning when such obstacles appear.
Overall, our results highlight PHANS as a scalable and practical approach for real-world applications such as high-density warehouse logistics and valet parking systems. Future research will focus on adaptive target-path planning and null-agent assignment strategies. Additionally, we will extend PHANS to handle more complex scenarios and arbitrary graphs.